### v1

# GeoEdit: Geometric Knowledge Editing for Large Language Models

Abstract
Regular updates are essential for maintaining up-to-date knowledge in large language models (LLMs). Consequently, various model editing methods have been developed to update specific knowledge within LLMs. However, training-based approaches often struggle to effectively incorporate new knowledge while preserving unrelated general knowledge. To address this challenge, we propose a novel framework called Geometric Knowledge Editing (GeoEdit). GeoEdit utilizes the geometric relationships of parameter updates from fine-tuning to differentiate between neurons associated with new knowledge updates and those related to general knowledge perturbations. By employing a direction-aware knowledge identification method, we avoid updating neurons with directions approximately orthogonal to existing knowledge, thus preserving the model’s generalization ability. For the remaining neurons, we integrate both old and new knowledge for aligned directions and apply a “forget-then-learn” editing strategy for opposite directions. Additionally, we introduce an importance-guided task vector fusion technique that filters out redundant information and provides adaptive neuron-level weighting, further enhancing model editing performance. Extensive experiments on two publicly available datasets demonstrate the superiority of GeoEdit over existing state-of-the-art methods.
1 Introduction
Figure 1: Conceptual illustration of F-Learning Ni et al. (2024) and our proposed GeoEdit.
Large language models (LLMs) have demonstrated the ability to store vast amounts of knowledge during pre-training and retrieve it during inference Yao et al. (2023); Wang et al. (2024e). However, much of the knowledge in the real world is constantly evolving. For instance, the answer to the question “Who is the President of the United States?” was “Joe Biden” in 2024, but it is now “Donald Trump”. As a result, some knowledge that was once correct in LLMs can become obsolete or inaccurate Li and Chu (2024); Huang et al. (2024).
To address this issue, model editing methods have been developed to update the target new knowledge while preserving unrelated general knowledge within the model Hong and Lipani (2024); Ma et al. (2024); Wang et al. (2024f). Specifically, current model editing methods typically follow the “locate-and-edit” paradigm Wang et al. (2023, 2024a); Li et al. (2024). The core idea is first to locate influential weights in LLMs and then edit them by introducing a perturbation. Although effective, these approaches incur significant computational overhead to identify the important neurons and parameters Meng et al. (2022). Some methods also require sampling additional data (e.g., from Wikipedia) to mitigate the impact on the general knowledge within LLMs during editing Meng et al. (2023); Fang et al. (2024), introducing extra costs and potential biases.
In contrast, fine-tuning with updated knowledge, as demonstrated in recent studies Zhao et al. (2024); Wang and Li (2024); Liu et al. (2024), offers a more straightforward solution through the use of parameter-efficient fine-tuning (PEFT) techniques. These methods employ various strategies to edit the model without the need to differentiate the importance of individual parameters Feng et al. (2025). Among these, the pioneering work F-Learning Ni et al. (2024) introduces a new learning framework called “Forgetting before Learning,” as illustrated in Figure 1(a). This approach is based on the empirical observation that new knowledge can be difficult to learn when it conflicts with existing knowledge. By first forgetting outdated knowledge, the learning of new knowledge becomes easier. However, this approach has a critical limitation: it struggles to balance the integration of new knowledge with the preservation of existing general knowledge. Specifically, F-Learning assumes that all updates between old and new knowledge are inherently conflicting, which oversimplifies the complexity of knowledge integration. Furthermore, the uncontrained forgetting process can significantly impact the model’s generalization ability to out-of-scope samples, considerably reducing the performance of the Locality metric (see Section 3).
To address these limitations, we propose Geometric Knowledge Editing (GeoEdit), a novel fine-tuning-based model editing framework that enhances editing precision while strongly preserving model generalization without the need for additional unrelated data. The core insight of GeoEdit is to distinguish between neurons associated with new knowledge updates and those linked to general knowledge perturbations by analyzing the geometric relationships of parameter updates caused by fine-tuning. By masking the updates of general-knowledge-related neurons, we prevent negative impacts on the model’s generalization ability. At the same time, we optimize the update strategy for new-knowledge-related neurons, further enhancing the effectiveness of model editing.
Specifically, we first fine-tune the current model separately on the old and new knowledge datasets. This allows us to derive neuron-level task vectors, $\tau_{old}$ and $\tau_{new}$, using task arithmetic Ilharco et al. (2022), which capture the directions of knowledge retention and updating w.r.t. each neuron, as shown in Figure 1(b). We then introduce a direction-aware knowledge identification method that computes the angle $\phi$ between these two directions to classify neurons, followed by customized editing strategies: (i) Orthogonal Knowledge Editing (for approximately orthogonal directions): Neurons with updates orthogonal to old knowledge are classified as general-knowledge-related neurons. These updates are considered detrimental to the model’s generalization ability, so we refrain from updating these neurons. The remaining neurons are treated as new-knowledge-related neurons, which are updated with two different strategies: (ii) Synergistic Knowledge Editing (for aligned directions): When there is slight conflict between old and new knowledge, we can leverage their similarities to simultaneously integrate both. (iii) Conflicting Knowledge Editing (for opposite directions): For updates with significant conflict, we apply the F-Learning strategy, where old knowledge is first forgotten before integrating new information.
Additionally, to mitigate angular bias in high-dimensional space, GeoEdit projects the task vectors into a lower-dimensional semantic space using an auto-encoder. Angle calculations and editing operations are performed in this compressed space, after which the decoder reconstructs the final editing task vector $\tau_{edit}$ for parameter updates. To optimize vector fusion, we introduce an importance-guided task vector fusion technique, which applies fine-grained weights to the vectors and suppresses noise from redundant parameters, further enhancing the effectiveness of model editing.
Our main contributions are summarized as:
  • We propose a novel geometric knowledge editing framework (GeoEdit) for updating LLMs.
  • We develop new direction-aware knowledge identification and importance-guided task vector fusion techniques.
  • Extensive evaluation on two widely-used datasets shows that GeoEdit overcomes the limitations of F-Learning, improving the Locality metric by 7.4% while maintaining the best performance in the Reliability and Generality metrics.
2 Related Work
Knowledge editing has gained significant attention due to the increasing need to update the knowledge in LLMs Wang et al. (2024e); Bi et al. (2024).
Existing methods can be classified as follows: (1) Fine-tuning is an intuitive and straightforward way to update the model’s knowledge Feng et al. (2023); Gangadhar and Stratos (2024); Zheng et al. (2024). Recent parameter-efficient fine-tuning methods like Prefix-Tuning Li and Liang (2021) and LoRA Hu et al. (2021) have made fine-tuning more suitable for knowledge editing. Ni et al. (2024) proposes the “Forgetting before Learning” (F-Learning) paradigm, which employs parametric arithmetic to facilitate the forgetting of old knowledge and the learning of new knowledge. However, this often leads to unintended changes in unrelated knowledge, negatively affecting generalization Wang et al. (2024b); Gu et al. (2024). (2) Meta-learning (“learning to learn”) methods use a hypernetwork to generate task-specific updates, allowing LLMs to quickly adapt to new data without retraining from scratch De Cao et al. (2021); Shi et al. (2024); Wang et al. (2024d). (3) Locate-and-edit methods usually locate influential parameters and then edit them by introducing a perturbation Zhang et al. (2024); Jiang et al. (2024); Xu et al. (2024). Classic methods like ROME Meng et al. (2022) and MEMIT Meng et al. (2023) use causal reasoning to identify key neuron activations and adjust specific weights. More recent approaches, such as AlphaEdit Fang et al. (2024), improve the method by projecting perturbations onto the null space of preserved knowledge.
However, fine-tuning and locate-and-edit methods struggle to balance new knowledge updates with preserving unrelated knowledge Gupta et al. (2024); Feng et al. (2024a). For instance, locate-and-edit methods typically require large additional datasets to capture general knowledge and avoid disruption during editing Wang et al. (2024c); Hsueh et al. (2024). However, these approaches introduce extra costs and biases.
Thus, our paper focuses on improving fine-tuning methods. Motivated by F-Learning Ni et al. (2024), we propose distinguishing between general knowledge and updated knowledge based on the angular divergence between the updated directions of old and new knowledge. This allows us to avoid updating general knowledge, ensuring model generalization while applying tailored strategies to improve the effectiveness of knowledge updates.
3 Preliminary
Figure 2: Overview of GeoEdit. Step (a): Neuron-level task vectors $\tau_{old}$ and $\tau_{new}$ are extracted for both the old and new knowledge datasets using parametric arithmetic. Step (b): An auto-encoder is trained to project a low-dimensional representation of the task vectors, eliminating the angular bias issue in high-dimensional space. Step (c): The latent task vectors, $h_{new}$ and $h_{old}$, are used to classify neurons based on the angle between their corresponding task vectors, followed by the application of different editing strategies to obtain $h_{edit}$. Finally, $h_{edit}$ is passed through the decoder to produce $\tau_{edit}$, which is added to the initial model to generate the edited model $f_{\theta_{e}}$.
Task Definition
Model editing, also referred to as knowledge updating, involves modifying the behavior of an initial target model on specific edit examples without compromising its performance on unrelated examples. More precisely, given an initial model $f_{\theta}$ and a set of input-output knowledge pairs $D_{old}=\{(x_{1},y_{1}),(x_{2},y_{2}),\dots,(x_{k},y_{k})\}$, the task is to update the model parameters to obtain a new model $f_{\theta_{e}}$ and a corresponding set of new input-output pairs $D_{new}=\{(x_{1},y_{1}^{\text{new}}),(x_{2},y_{2}^{\text{new}}),\dots,(x_{k},y_{k}^{\text{new}})\}$, where $k$ denotes the number of knowledge pairs to be updated. The objective of the post-edit model $f_{\theta_{e}}$ is to meet three essential properties: reliability, generality, and locality Wang et al. (2024b).
Reliability
Reliability measures the accuracy of the updated model on the new knowledge. Specifically, the output for “Who is the President of the US?” should be updated from “Joe Biden” to “Donald Trump.” This can be formalized as follows:
$\mathbb{E}_{x_{e},y_{e}\sim D_{new}}\mathbbm{1}\left\{\text{argmax}_{y}f_{\theta_{e}}(y|x_{e})=y_{e}\right\}.$ (1)
Generality
Generality means that the new model $f_{\theta_{e}}$ should also update rephrased in-scope examples $I(x_{e},y_{e})$. Such as the answer to “Who holds the position of the President of the US?” should also be changed from “Joe Biden” to “Donald Trump”. This is evaluated by the average accuracy of $f_{\theta^{*}}$ on examples from the equivalence neighborhood, as expressed by:
$\mathbb{E}_{x^{\prime}_{e},y^{\prime}_{e}\sim I(x_{e},y_{e})}\mathbbm{1}\left\{\text{argmax}_{y}f_{\theta_{e}}(y|x^{\prime}_{e})=y^{\prime}_{e}\right\}.$ (2)
Locality
A good edit should modify relevant knowledge without affecting other irrelevant out-of-scope examples $O(x_{e},y_{e})$. For example, the question, “Who said: this is a battle for the soul of the nation?” should remain unchanged as “Joe Biden”. Locality (or specificity) is defined as:
$\mathbb{E}_{x^{\prime}_{e},y^{\prime}_{e}\sim O(x_{e},y_{e})}\mathbbm{1}\left\{\text{argmax}_{y}f_{\theta_{e}}(y|x^{\prime}_{e})=f_{\theta}(y|x^{\prime}_{e})\right\}.$ (3)
4 Proposed Method: GeoEdit
In this section, we present our method for knowledge editing in LLMs. As illustrated in Figure 2, GeoEdit follows a three-step process:
4.1 Extracting Neuron-level Task Vectors
Supervised fine-tuning (SFT) on a dataset injects new knowledge into the LLMs, reflected in model parameter changes. For an initial model $f_{\theta}$ with parameters $\theta$, fine-tuning on dataset $D$ produces updated parameters. The difference between the updated and original parameters is referred to as the task vector Ilharco et al. (2022), calculated as:
$\tau=\mathrm{FT}\{\theta,D\}-\theta$ (4)
where $\tau$ is the corresponding task vector, and FT is the fine-tuning operation. Unlike F-Learning, we fine-tune the initial model $f_{\theta}$ separately on the old and new knowledge datasets to isolate their respective adaptations, then compute task vectors:
$\tau_{old}=\mathrm{FT}\{\theta,D_{old}\}-\theta$ (5)
$\tau_{new}=\mathrm{FT}\{\theta,D_{new}\}-\theta$ (6)
where $D_{old}$ and $D_{new}$ are datasets encoding outdated and updated knowledge respectively.
While prior research typically captures task vectors at the model level Ilharco et al. (2022), we propose extracting them at the neuron level for finer control. Let $\theta=\{\theta^{1},\theta^{2},\dots,\theta^{N}\}$ represent the $N$ neurons in the LLM, where the $i$-th neuron is represented by $\theta^{i}\in\mathbb{R}^{d_{n}}$ with $d_{n}$ dimensional parameters. The neuron-level task vectors are given by $\tau_{new}=\{\tau_{new}^{1},\tau_{new}^{2},\dots,\tau_{new}^{N}\}$, where $\tau_{new}^{i}$ corresponds to the new knowledge task vector for the $i$-th neuron. This approach enables more granular analysis of parameter changes and selective editing of knowledge-specific neurons, enhancing model editing precision.
After obtaining the task vectors for both old and new knowledge, we focus on the directional characteristics, which are more crucial than magnitudes for knowledge editing. We define the direction of $\tau_{old}$ as the knowledge retention direction and $\tau_{new}$ as the knowledge updating direction. By analyzing the angle between these directions, we can distinguish general-knowledge-related neurons to avoid harming generalization and new-knowledge-related neurons to enhance editing effectiveness.
4.2 Dimensionality Reduction Using AE
After obtaining task vectors $\tau_{old}$ and $\tau_{new}$, we use an auto-encoder (AE) to map them into a low-dimensional latent space for knowledge editing. This step mitigates the angular bias inherent in high-dimensional task vectors. For example, in LLaMA-7B, each neuron-level task vector has 4096 dimensions, which naturally tend toward orthogonality, distorting the interpretation of their relationships. Dimensionality reduction also removes irrelevant information, improving the editing accuracy.
Thus, we train a semantic encoder and decoder, both implemented using multi-layer perceptrons (MLPs). The encoder maps high-dimensional task vectors to a lower-dimensional space for editing, while the decoder reconstructs the edited vectors, which are then added back to the initial model for the final results. Specifically, the semantic encoder, denoted as SemEnc($\cdot$), maps the high-dimensional task vectors $\tau_{old}$ and $\tau_{new}$ into the latent space as:
$h^{i}=\text{SemEnc}(\tau^{i})$ (7)
where $h^{i}\in\mathbb{R}^{d_{latent}}$ is the latent task vector, and $d_{latent}$ denotes its dimensionality. The decoder, Dec($\cdot$), then generates $\hat{\tau}^{i}$ from $h^{i}$ as follows:
$\hat{\tau}^{i}=\text{Dec}(h^{i})$ (8)
where $\hat{\tau}^{i}$ is the reconstructed task vector. The auto-encoder is optimized using both a reconstruction loss and a semantic consistency loss:
[…]
where MSE($\cdot$) is mean square error loss function, and KL($\cdot$) is the Kullback-Leibler divergence.
4.3 Geometric Knowledge Editing
After training the auto-encoder, we project $\tau_{old}$ and $\tau_{new}$ into the low-dimensional latent space for knowledge editing. GeoEdit then edit the latent task vectors to obtain $h_{edit}$, which is decoded to produce the edited task vector $\tau_{edit}$. This vector is subsequently added to the initial model, resulting in the final edited model $f_{\theta_{e}}$.
Direction-aware Knowledge Identification
For neuron $i$, we first use the encoder to reduce the dimensionality of $\tau_{old}^{i}$ and $\tau_{new}^{i}$ yielding the latent task vectors $h_{old}^{i}$ and $h_{new}^{i}$. We then compute the angular divergence $\phi$ between $h_{old}^{i}$ and $h_{new}^{i}$ as:
$\phi=\arccos\frac{h_{old}^{i}\cdot h_{new}^{i}}{\left|h_{old}^{i}\right|\cdot\left|h_{new}^{i}\right|}$ (10)
Neurons with angles near orthogonality (within the range of $\phi_{1}$ to $\phi_{2}$) are classified as general-knowledge-related, while the remaining neurons are classified as new-knowledge-related.
Importance-guided Task Vector Fusion
We then apply customized editing strategies based on the classification of neurons as follows:
$h_{edit}^{i}=\begin{cases}\alpha^{i}h_{old}^{i}+\beta^{i}h_{new}^{i},&\text{if$\phi\in(0^{\circ},\phi_{1})$}\\ 0,&\text{if $\phi\in[\phi_{1},\phi_{2}]$}\\ -\alpha^{i}h_{old}^{i}+\beta^{i}h_{new}^{i},&\text{if $\phi\in(\phi_{2},180^{\circ})$}\\ \end{cases}$ (11)
where $\alpha^{i},\beta^{i}\in[0,1]$ are the fusion weights, automatically assigned based on the neuron’s importance to both new and old knowledge, removing the need for manual adjustment.
To calculate the fusion weights, we measure the importance of each neuron by analyzing the gradient trajectory of its parameters during fine-tuning. The importance is determined by the collective contribution of its trainable parameters:
$\mathcal{I}(\theta^{i})=\frac{1}{d_{n}}\sum\limits_{j=1}^{d_{n}}s(w_{j})$ (12)
where $w_{j}$ represents the trainable parameters and $d_{n}$ is the total number of parameters in neuron $\theta^{i}$. The function $\mathcal{I}(\theta^{i})$ reflects the importance of the neuron, with higher values indicating greater significance. The function $s(\cdot)$ computes the importance of individual parameters based on the magnitude of the gradient-weight product Zhang et al. (2023):
$s\left(w\right)=\left|w\nabla_{w}\mathcal{L}\right|$ (13)
Due to stochastic sampling and training dynamics, the metric in Eq. (13) may vary, reducing reliability Feng et al. (2024b). To address this, we apply an exponential moving average to smooth the trajectory gradients across training iterations.
We normalize the importance scores $\mathcal{I}_{\text{old}}=\{\mathcal{I}_{\text{old}}^{1},\dots,\mathcal{I}_{\text{old}}^{N}\}$ and $\mathcal{I}_{\text{new}}=\{\mathcal{I}_{\text{new}}^{1},\dots,\mathcal{I}_{\text{new}}^{N}\}$ independently to the range [0, 1]. This yields the final fusion weights $\alpha=\{\alpha^{1},\dots,\alpha^{N}\}$ and $\beta=\{\beta^{1},\dots,\beta^{N}\}$, which are then applied to the corresponding task vectors $h_{\text{old}}$ and $h_{\text{new}}$ for editing.
By applying Eq. (11), our GeoEdit effectively addresses the challenges in model editing:
  • Preserving general knowledge (Case 2): We mask updates to general-knowledge-related neurons to avoid negatively impacting the model’s generalization ability.
  • Improving knowledge editing (Case 1 &\&& 3): For acute angles, we leverage the similarity between old and new knowledge for efficient integration. For obtuse angles, significant conflict triggers a “forget-then-learn” strategy, optimizing the updates for new-knowledge-related neurons.
5 Experiments and Analysis
Dataset Method LLAMA2-7B LLAMA-7B
Reliability Generality Locality Reliability Generality Locality
ZsRE Original model 43.70 43.17 / 43.29 42.85 /
LoRA 43.10 42.20 70.83 46.93 45.87 75.86
F-LearningLoRA 46.91 46.21 72.50 47.56 46.90 76.09
GeoEditLoRA 46.11 47.19 78.42 48.56 47.11 80.50
FT-c 49.02 46.96 67.37 47.33 45.51 68.14
Full-FT 81.02 74.67 70.51 70.52 66.69 65.26
MEND 29.77 25.86 71.54 30.99 27.12 69.83
ROME 43.67 42.66 93.14 43.45 42.94 98.60
MEMIT 83.57 79.06 70.52 78.30 77.43 69.44
RECT 84.08 77.80 69.03 78.78 76.20 67.97
F-LearningFT 84.65 81.51 70.92 83.06 79.50 70.09
GeoEditFT 85.21 82.43 75.71 84.81 79.86 75.15
CounterFact Original model 18.47 16.95 / 21.61 17.88 /
LoRA 30.56 23.24 40.08 27.54 21.21 39.75
F-LearningLoRA 31.17 23.63 40.42 29.47 22.89 44.91
GeoEditLoRA 31.64 24.13 43.83 30.89 23.00 47.68
FT-c 29.23 19.32 19.70 26.97 17.90 20.09
Full-FT 65.99 44.08 28.34 32.13 31.95 32.51
MEND 14.77 14.67 90.93 17.51 16.27 89.64
ROME 18.41 17.20 93.60 21.83 19.08 92.27
MEMIT 61.94 37.45 21.90 56.94 31.48 25.70
RECT 62.90 39.86 20.03 57.82 33.51 23.48
F-LearningFT 69.53 45.56 28.41 56.39 39.75 31.87
GeoEditFT 68.34 46.53 37.73 55.88 40.60 42.33
Table 1: Results on three metrics for the two datasets using LLAMA2-7B and LLAMA-7B.
Datasets
We use two widely recognized datasets: ZsRE Levy et al. (2017) and CounterFact (Meng et al., 2022). ZsRE is a Question Answering dataset that generates equivalence neighborhoods using question paraphrases produced through back-translation. CounterFact presents a more complex challenge with counterfactual data. We adopt the experimental setup from Yao et al. (2023), using the eval and edit subsets consisting of 19,085 and 10,000 examples, respectively. The datasets are partitioned into old and new knowledge categories, as in F-Learning Ni et al. (2024). For example, in ZsRE, old knowledge is modified to new knowledge, such as the change from “Los Angeles” to “New Orleans.” Further details and additional examples are provided in Appendix A.
Baselines
We evaluate GeoEdit against two types of methods: fine-tuning-based approaches, including full fine-tuning (Full-FT), LoRA Hu et al. (2021), FT-c Zhu et al. (2020), and F-Learning Ni et al. (2024); and locate-and-edit-based methods, including MEND Mitchell et al. (2022), ROME Meng et al. (2022), MEMIT Meng et al. (2023), and RECT Gu et al. (2024). Detailed descriptions are provided in the Appendix B.
Training Details
Following the setup of F-Learning Ni et al. (2024), we use LLAMA2-7B and LLAMA-7B as the base models and focus on updating old knowledge to new knowledge. First, we fine-tune the base model on the old knowledge for three epochs using full fine-tuning, resulting in the original model, which serves as the baseline for our experiments. At this stage, the model has fully learned the old knowledge, making it essential to analyze the angular divergence between the task vectors of the old and new knowledge. In GeoEdit, the encoder and decoder consists of 2-layer MLPs with dimensions $[4096\rightarrow 2048,2048\rightarrow 512]$ and $[512\rightarrow 2048,2048\rightarrow 4096]$, respectively, where $d_{latent}$ is set to 512. We set $\lambda$ in Eq. (9) to 0.5, $\phi_{1}$ and $\phi_{2}$ in Eq. (11) to $85^{\circ}$ and $95^{\circ}$, respectively. Further details on the experimental setup can be found in Appendix C.
5.1 Experimental Results
The overall results are presented in Table 1. Our GeoEdit method yields promising results in LoRA and full fine-tuning, surpassing other baselines. Surprisingly, ROME maintains Reliability and Generality almost unchanged from the original model across both datasets while achieving high Locality (greater than 90). Since the injection of new knowledge typically impacts Locality, this suggests that ROME performs minimal knowledge updating, likely due to its limited parameter edits. Although F-Learning achieves high Reliability and Generality, its Locality performance drops significantly due to the lack of constraints during the forgetting phase, adversely harms the model’s generalization ability. In contrast, our GeoEdit method, which updates parameters based on the direction of new and old knowledge, improves Locality by 7.4% over F-Learning by masking orthogonal updates. Additionally, by classifying different knowledge editing strategies for new-knowledge-related neurons, our method further improves Reliability and Generality. These results demonstrate that GeoEdit achieves a better balance between updating new knowledge and preserving general knowledge.
Method Reliability Generality Locality
GeoEdit 85.21 82.43 75.71
- Synergistic 84.37 81.13 75.94
- Orthogonal 85.29 82.86 71.70
- Conflict 82.57 79.40 75.45
+ MW 84.91 82.04 73.73
Table 2: Ablation study. “- Synergistic”, “- Orthogonal”, and “- Conflict” refer to removing the synergistic, orthogonal, and conflict knowledge editing strategies, respectively. “+ MW” denotes replacing the importance-guided fusion with a manually set weighting approach.
5.2 Ablation Study
We conduct ablation studies to evaluate the effectiveness of the techniques in GeoEdit. The results on the ZsRE dataset with LLaMA2-7B are shown in Table 2. Additional analysis of hyperparameter sensitivity is provided in Appendix D.
Effect of Geometric Editing Strategies.
In GeoEdit, we categorize knowledge updates into three strategies based on the angle between the directions of knowledge retention and knowledge updating: synergistic, orthogonal, and conflict knowledge editing. To evaluate their impact, we disable each strategy and replace it with vanilla fine-tuning on new knowledge. For example, “- Orthogonal” means setting $h_{edit}=h_{new}$ instead of $h_{edit}=0$.
As shown in Table 2, removing any of the editing strategies leads to performance degradation. Notably, excluding the orthogonal knowledge editing strategy significantly reduces locality, from 75.7% to 71.7%, highlighting the close relationship between orthogonal parameter updates and the model’s general knowledge. Moreover, removing the conflict knowledge editing most adversely affects the reliability metric, causing a decline from 85.2% to 82.6%. These results emphasize the importance of each editing strategy.
$h_{latent}$ Reliability Generality Locality
128 44.94 43.79 75.68
256 46.75 46.54 77.33
512 46.11 47.19 78.42
1024 25.39 24.49 94.81
2048 23.03 24.14 96.14
Table 3: Ablation study on latent dimension $h_{\text{latent}}$.
Figure 3: Distribution of the angles $\phi$ between task vectors before and after dimensionality reduction.
Figure 4: Visualization of the magnitudes of task vectors $\tau_{\text{old}}$ and $\tau_{\text{new}}$ along with the importance-guided fusion weights. All results are normalized to the range of 0 to 1.
Effect of Importance-guided Task Vector Fusion.
We replace the importance-based weights $\alpha$ and $\beta$ in Eq. (11) with manually set values (“+ MW”), applying the same weight to all neurons instead of assigning neuron-specific weights as in GeoEdit. Through grid search, we set $\alpha=0.3$ and $\beta=1$.
The performance decline in Table 2 illustrates the effectiveness of our importance-guided fusion mechanism. This approach offers two key advantages: First, it provides neuron-level adaptive weights, ensuring greater precision than manually set strategies. Second, it ensures that parameter updates are influenced not only by the task vector’s magnitude but also by each neuron’s importance. The smaller weight “masks” significant parameter changes for less important neurons, minimizing their impact on the model’s generalizability.
Effect of Latent Space Dimension.
We investigate the effect of the latent dimension $h_{latent}$ in the auto-encoder on model performance. As shown in Table 3, larger latent dimensions (greater than 1024) cause performance collapse, suggesting that they still suffer from angular bias, with task vectors tending toward orthogonality. Conversely, smaller latent dimensions (less than 256) result in excessive dimensionality reduction, hindering effective training of the auto-encoder and leading to performance degradation. Our experiments demonstrate that a latent dimension of 512 strikes an optimal balance, capturing sufficient information while avoiding unnecessary complexity, which ensures effective model editing and strong generalization.
5.3 Visualization
We present two key visualizations to demonstrate the effectiveness of our approach:
Angle Distribution Between Old Knowledge Retention and New Knowledge Updating Directions.
We visualize the distribution of the angle $\phi$ between task vectors before and after dimensionality reduction using the auto-encoder, as shown in Figure 3. In the original high-dimensional space, the angles are predominantly concentrated around 90 degrees, as high-dimensional vectors tend to be nearly orthogonal. To mitigate this issue, we trained an auto-encoder to reduce the dimensionality of the task vectors. After dimensionality reduction, the angle distribution spans the full range from 0 to 180 degrees, revealing various types of conflicts between old and new knowledge. This motivates the development of different editing strategies based on angles, allowing us to distinguish between updates that correspond to learning new knowledge and those that modify general knowledge. Thus, we propose a direction-aware geometric model editing approach that effectively handles distinct knowledge updates to enhance model editing.
Visualization of Task Vector Magnitudes and Importance-guided Fusion Weights.
Figure 4 illustrates that while the magnitudes of the task vectors $\tau_{\text{old}}$ and $\tau_{\text{new}}$ are generally large, only a subset of the parameters are truly important, highlighting redundancy in the task vectors. Our importance-guided fusion mechanism effectively filters out this redundancy, enhancing the model editing process and minimizing its impact on generalization.
5.4 Editing Time Analysis
To evaluate the efficiency of our method, we present the average time required to edit 1000 samples in Table 4. We observe that fine-tuning-based methods are significantly faster than location-based methods, as the latter rely heavily on the specific positions of neurons and parameters, increasing complexity and editing time. Among fine-tuning approaches, F-Learning, which follows a two-stage process of forgetting before learning, takes approximately twice as long as Full-FT. In contrast, our method enables the parallel acquisition of old and new knowledge, resulting in training times comparable to Full-FT. The additional time primarily comes from training the auto-encoder (about 2.5 minutes) and editing the task vector (around 68 seconds). Thus, our method requires less time than F-Learning, delivering substantial performance improvements. This advantage becomes more pronounced as the number of samples to edit increases.
Method Average time per 1000 edits
zsRE COUNTERFACT
FT-c 653.2(s) 579.3(s)
ROME 8736.6(s) 7241.7(s)
MEMIT 3448.8(s) 3390.8(s)
Full-FT 810.2(s) 792.4(s)
F-LearningFT 1670.4(s) 1603.8(s)
GeoEditFT 1028.0(s) 1010.6(s)
Table 4: Editing time for two datasets on LLAMA2-7B.
6 Conclusion
In this paper, we introduce Geometric Knowledge Editing (GeoEdit), a novel framework that utilizes the geometric relationships between parameter updates to improve model editing. By applying a direction-aware knowledge identification technique, GeoEdit classifies neurons into two categories: general-knowledge-related neurons, whose parameter updates are masked to prevent negative impacts on model generalization, and new-knowledge-related neurons, where an importance-guided task vector fusion technique is applied to enhance editing. Extensive experiments demonstrate the effectiveness of GeoEdit for model editing.
Limitations
We acknowledge two limitations in this work.
First, GeoEdit requires access to old knowledge datasets to extract the task vector $\tau_{old}$. In some cases, however, such datasets may not be available, meaning we only know the updated results. A potential solution is to input the task to be edited directly into the initial model and use the output as the old knowledge. However, this introduces additional inference costs, especially in our mass-editing settings. Furthermore, for open-ended questions, selecting the appropriate output as the reference is another challenge. We plan to explore ways to extend GeoEdit to address these issues and improve its adaptability.
Second, the core of GeoEdit relies on using the angle between parameter updates to differentiate between disturbances to general knowledge and the learning of new knowledge. While this approach offers valuable insights, it still results in some loss of model generalization, suggesting that the angle alone cannot fully decouple new knowledge learning from general knowledge disturbance. To address this, we aim to consider multiple geometric variables, such as task vector projections and magnitude, to further refine GeoEdit and enhance performance in the future.
Appendix A Datasets and Examples
We follow the F-Learning approach Ni et al. (2024), which divides datasets into old and new knowledge. Below, we provide an overview of the datasets used, with detailed descriptions available in the original F-Learning paper. We use two well-known datasets: ZsRE Levy et al. (2017) and CounterFact (Meng et al., 2022). ZsRE is a Question Answering (QA) dataset that incorporates question rephrasings via back-translation Lu et al. (2021a); Feng et al. (2024c), while CounterFact is a more challenging counterfactual dataset. We use the eval and edit sets, containing 19,085 and 10,000 examples, respectively. Here’s an example from the ZsRE dataset:
{"subject": "Watts Humphrey", "src": "What university did Watts Humphrey attend?", "pred": "Trinity College", "rephrase": "What university did Watts Humphrey take part in?", "alt": "University of Michigan", "answers": ["Illinois Institute of Technology"], "loc": "nq question: who played desmond doss father in hacksaw ridge", "loc-ans": "Hugo Weaving", "cond": "Trinity College >> University of Michigan || What university did Watts Humphrey attend?"}
In this example, old knowledge ("Trinity College") is replaced with new knowledge ("University of Michigan") for the same question. The "rephrase" field evaluates the model’s generalization, while "loc" assesses the locality of the model’s output. The datasets are divided into old and new knowledge, with the same format maintained for effective supervised fine-tuning. Below are examples of old and new knowledge in an instruction-based format:
Old knowledge:
{"instruction": "What university did Watts Humphrey attend?", "input": "", "output": "Trinity College" }
New knowledge:
{"instruction": "What university did Watts Humphrey attend?", "input": "", "output": "University of Michigan" }
It’s important to note that old knowledge represents correct real-world facts, while new knowledge is deliberately incorrect, ensuring that the original model has not previously learned it. This setup avoids ambiguity in determining whether the new knowledge was already part of the model’s prior knowledge.
$\lambda$ Reliability Generality Locality
0 46.07 46.81 76.65
0.3 46.21 47.59 77.74
0.5 46.11 47.19 78.42
0.7 46.25 47.70 77.41
0.9 46.40 46.87 77.06
Table 5: Performance comparisons of GeoEdit equipped with different $\lambda$.
$\phi_{1}$ $\phi_{2}$ Reliability Generality Locality
$87^{\circ}$ $93^{\circ}$ 48.07 49.13 74.81
$75^{\circ}$ $105^{\circ}$ 44.57 44.81 83.63
$80^{\circ}$ $100^{\circ}$ 45.86 46.55 80.26
$85^{\circ}$ $95^{\circ}$ 46.11 47.19 78.42
Table 6: Performance comparisons of GeoEdit equipped with different $\phi$.
Appendix B Baseline Details
We evaluate our GeoEdit method against a range of fine-tuning and locate-and-edit-based approaches.
For fine-tuning methods, we first compare our approach with full fine-tuning (Full-FT) and LoRA Hu et al. (2021). LoRA (Low-Rank Adaptation) introduces small, trainable matrices into each layer of the model, enabling efficient adaptation while keeping most of the pre-trained parameters frozen. We also evaluate FT-c Lu et al. (2021b), a fine-tuning method that applies an $L_{\infty}$ constraint to help retain irrelevant knowledge. Additionally, we compare with the F-Learning method Ni et al. (2024), which first forgets outdated knowledge to facilitate the incorporation of new information.
For locate-and-edit-based methods, we start by evaluating MEND Mitchell et al. (2022), which learns a hypernetwork to generate weight updates by decomposing fine-tuning gradients. We also experiment with ROME Meng et al. (2022), a method that updates specific factual associations through causal intervention. Additionally, we compare with MEMIT Liu et al. (2023), a method designed for directly updating large-scale memories. Finally, we include RECT Gu et al. (2024), which regularizes edit updates by imposing constraints on the complexity of the weight changes.
Appendix C Implementation Details of Experiments
Here we will introduce more completion details and settings of experiments. First, we used LLAMA2-7B and LLAMA-7B as the base models, and then we trained the base model on the old knowledge for 3 epochs by full fine-tuning to simulate an original model that has fully learned old knowledge for our experiments. This makes the forgetting operation more reasonable and effective, and at the same time tries to avoid the problem of being unable to determine whether the new knowledge output by the LLM is learned from the data or commanded by itself as mentioned above. During testing, we use a greedy decoding strategy to ensure the uniqueness of the model’s output. All experiments were conducted on a setup using 4 × A100-80G GPUs.
For full-parameter fine-tuning, all experiments were conducted for 3 epochs with a batch size of 4, gradient accumulation steps set to 4, and a learning rate of 5e-5. Specifically, for LoRA fine-tuning, the hyperparameters were set as follows: $r=8$, $\alpha=32$, dropout = 0.05, with the targeting modules being [q_proj, v_proj]. When we used the Deepspeed, we set 4 processes and zero-stage is 2. The experiment settings for the baseline methods, including F-Learning, FT-c, and ROME, follow those outlined in Ni et al. (2024).
It is worth noting that we used the same hyperparameters across different datasets and backbones, demonstrating the generalizability of our method without requiring extensive hyperparameter tuning for each specific setting.
Appendix D Sensitivity Analysis for Hyperparameters
The proposed framework incorporates two key hyperparameters: $\lambda$, which balances the autoencoder loss in Eq. (9), and $\phi$, which defines the thresholds for dividing different editing strategies. Our analysis aims to assess the impact of varying these hyperparameters on the performance of our method, with tests conducted on the ZsRE dataset using LLaMA2-7B backbone model (LoRA fine-tuning).
As shown in Table 5, we determine that the optimal setting for $\lambda$ is 0.5. Regarding the selection of the threshold for dividing editing strategies, the article sets $\phi_{1}$ and $\phi_{2}$ to $85^{\circ}$ and $95^{\circ}$, respectively. Table 6 below shows the model’s performance with varying thresholds for $\phi$. It can be seen that as the range between $\phi_{1}$ and $\phi_{2}$ increases, meaning more updates are masked, this better prevents interference with the model’s general knowledge but limits the learning of new knowledge. This results in an increase in locality but a decrease in reliability. Conversely, narrowing the range of $\phi_{1}$ and $\phi_{2}$ enhances the model’s ability to update, but it also impacts its generalization ability. Therefore, we choose the range of $85^{\circ}$ to $95^{\circ}$ as the optimal balance for masking, achieving the best trade-off between learning new knowledge and preserving general knowledge.